OpenStudy (anonymous):

how many milliliters of a 15% saline and how many milliliters of a 40% saline solution must be mixed to produce 10 milliliters of a 30% saline solution?

OpenStudy (anonymous):

x+y=10 15x+40y=10(30) Solve the system.

OpenStudy (anonymous):

would i be using the elimination method to solve?

OpenStudy (anonymous):

Yes, or substitution.

OpenStudy (anonymous):

thnks i got it!

OpenStudy (anonymous):

Can you help me with this one too please? A total of $9,600 was invested in two separate accounts earning 5.5% and 3.75% annual interest. If the total simple interest earned for the year was $491.25, then how much was invested in each account?

OpenStudy (anonymous):

x+y=9600 .055x+.0375y=491.25 This one is a little messier, but you can solve it the same way you did the other.

OpenStudy (anonymous):

thanks again @AnimalAin

OpenStudy (anonymous):

@Luis_Rivera i got x=4 and y =6

OpenStudy (anonymous):

No sweat. Do math every day.